I am not sure if this is trig but the question is
Tn = 2n + 3
write the first five terms of the sequence

Answer by Edwin McCravy(20086)

I am not sure if this is trig but the question is
Tn = 2n + 3
write the first five terms of the sequence

No it's not trig.

It means to replace n by 1, then by 2, then by 3, then by 4, then by 5.

Tn = 2n+3

Replacing n by 1:

T1 = 2(1)+3 = 2+3 = 5

Replacing n by 2:

T2 = 2(2)+3 = 4+3 = 7

Replacing n by 3:

T3 = 2(3)+3 = 6+3 = 9

Replacing n by 4:

T4 = 2(4)+3 = 8+3 = 11

Replacing n by 5:

T5 = 2(5)+3 = 10+3 = 13

and the sequence is

T1, T2, T3, T4, T5

or  5, 7, 9, 11, 14

Edwin